From the following data of Megh Enterprises.


a. Calculate the combined breakeven sales. The Company is producing three products.


product, sales, Variable cost

A 10000 6000

B 5000 2500

C 5000 2000

The company incurred a fixed cost of Rs 5700.


b. Also, Discuss the concept and relevance of Breakeven sales for an enterprise


1
Expert's answer
2022-04-01T16:35:34-0400

"Breakeven sales = \frac {fixed costs * sales}{sales - variable costs}"


Fixed costs = 5700

Sales = "10000+5000+5000 = 20000"

Variable costs = "6000+2500+2000 = 10500"

"Breakeven sales = \frac {{5700}\times{20000}}{20000 - 10500}"


"Breakeven sales = \frac{114000000}{9500}"


Breakeven sales = 12000